OpenStudy (anonymous):

algebra 2 help please

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

how to divide this

OpenStudy (dumbcow):

when dividing fractions ... multiply by reciprocal \[\frac{\frac{x^{2}-4}{x-8}}{x-2} = \frac{x^{2}-4}{x-8}*\frac{1}{x-2} = \frac{x^{2}-4}{(x-8)(x-2)}\]

OpenStudy (dumbcow):

now factor the top and see what cancels

OpenStudy (anonymous):

so is answer (x-2) (x+2 / ( x-8)?

OpenStudy (dumbcow):

almost ....the (x-2) cancels leaving \[\frac{x+2}{x-8}\]